A 900-kg car is going 20 m/s along a level road. How large a constant retarding (friction)

force is required to stop it at a distance of 30m? (hint: find the negative acceleration).


1
Expert's answer
2021-05-10T13:58:04-0400

"s=\frac{v^2-v_0^2}{2a}\to a=\frac{v^2-v_0^2}{2s}=\frac{0^2-20^2}{2\cdot30}=-6.67\ (m/s^2)"


"F=ma\to F=900\cdot(-6.67)=-6000\ (N)" . Answer